tinaa212:

Williamston Daily is a new daily newspaper. When they first started printing three years ago they had 5900 home delivery subscribers. The number of subscribers increases by 66 every month. Write an equation that expresses the number of subscribers B in terms of n, the number of months the paper has been printed.

Mercury:

model the quation after the generic equation y = mx + b, where y is the independent variable, m is the rate of change, x is the dependent variable, and b is the initial value of the independent variable start by identifying which variable is the independent variable, and which one is dependent variable (the # of subscribers, or the # of months? which one is which?)

Mercury:

answer choices, just so we have them on hand: a. B = 5900n + 66 c. B = 66n + 5900 b. B = 198n + 5900 d. B = 66n + 3